Suppose that the width of a certain rectangle is 5 inches less than its length. The area is numerically 16 less than twice the perimeter. Find the length and width of the rectangle.

Answer by mananth(16946)

Suppose that the width of a certain rectangle is 5 inches less than its length. The area is numerically 16 less than twice the perimeter. Find the length and width of the rectangle.
...
length = L
width = L-5
..
Area = L(L-5)
...
Perimeter = 2(L+L-5)
Perimeter = 2(2L-5)
...
L(L-5)=2*2(2L-5)-16
L^2-5L=8L-20-16
L^2-5L-8L=-20-16
L^2-13L=-36
L^2-13L+36=0
L^2-9L-4L+36=0
L(L-9)-4(L-9)=0
(L-9)(L-4)=0
So L= 4 OR 9
The dimensions are 9 by 4 feet